Glauber’s salt is a form of hydrated sodium sulfate, Na2SO4ᐧ xH2O(s), that is used medicinally as a mild laxative. Chemical analysis of this hydrate shows that it is 44.1% sodium sulfate. Determine the FORMULA and the NAME of this hydrate. 


1
Expert's answer
2021-04-25T23:11:43-0400

Q184524


Glauber’s salt is a form of hydrated sodium sulfate, Na2SO4ᐧ xH2O(s), that is used medicinally as a mild laxative. Chemical analysis of this hydrate shows that it is 44.1% sodium sulfate. Determine the FORMULA and the NAME of this hydrate. 


Solution :

If there is 44.1 % of sodium sulfate in the salt,

then the remaining (100 - 44.1 ) = 55.9 % will be H2O.


Let us consider that there are 100 grams of hydrated sodium sulfate.


Since there is 44.1 % sodium sulfate, so 100 grams of hydrated sodium sulfate will contain

44.1 grams of sodium sulfate and 55.9 grams of H2O.


Step 1: To find moles of Na2SO4 and H2O present in 100 grams of hydrated salt.


Convert 44.1 grams of Na2SO4 and 55.9 grams of H2O to moles by using their molar mass.


Molar mass of Na2SO4 = 142.041 g/mol


Molar mass of H2O = 18.0148 g/mol




"moles \ of \ Na_2SO_4 = 44.1 \ \bcancel{g \ Na_2SO_4 } * \frac{1 \ mol \ Na_2SO_4 }{142.041 \bcancel{g \ Na_2SO_4 }}"

"= 0.3105 \ mol \ of \ Na_ 2SO_4"



"moles \ of \ H_2O= 55.9 \ \bcancel{g \ H_2O } * \frac{1 \ mol \ H_2O}{18.0148 \bcancel{g \ H_2O }}"

"= 3.1030 \ mol \ of \ H_2O"


For finding the number 'x' . We will divide moles of both by the one with the least number of moles.

Here Na2SO4 has least number of moles.



"For Na_2SO_4 , \ \ \ \ \ \frac{0.3105 \ mol }{0.3105 \ mol } = 1"





"For H_2O , \ \ \ \ \ \ \frac{3.1030 \ mol }{0.3105 \ mol } = 9.994 \ ≈ \ 10"

So for each mole of Na2SO4, there is 10 mole of H2O.


So x = 10.


The formula of the hydrated salt will be Na2SO4.10 H2O.


There are 10 water molecules in this formula. For denoting the number of hydrated water molecules we will use the prefix ' deca' .


10 means deca.


So the name of the compound will be


Sodium sulfate decahydrate.